Hi, Im Beckah. Im having issues with my math homework. Im in geometry, but we have an algebra problem in our homework. The question is " Which equation represents the following problem? Fifteen minus three times a number(n) equels negitive twenty-two. Find the number." So, I found the equation - it's 15-3n=-22 , right? So, now Im trying to find the number, but everytime I solve for it I get the wrong number, I know this becuase I go back and check, but I don't get -22. Do you think you could help me out? Thank you. :)
Found 2 solutions by stanbon, nerdybill:
Answer by stanbon(75887)

Which equation represents the following problem? Fifteen minus three times a number(n) equels negitive twenty-two. Find the number." So, I found the equation - it's 15-3n=-22 , right?
----
Right.
15-3n= -22
3n = 15+22
3n = 37
n = 37/3 = 12 1/3
-----------
Checking:
15 - 3(37/3) = -22
15 - 37 = -22
-22 = -22
==================
Cheers,
Stan H.

Answer by nerdybill(7384)

Starting with:
15-3n=-22
we subtract 15 from both sides:
15-3n-15 = -22-15
-3n = -37
divide both sides by -3:
n = 37/3 (this is your answer)
.
To check, we substitute back in:
15-3n=-22
15-3(37/3)=-22
15-37=-22
-22 = -22